# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v1

@@ 2.2 RADI-based Non-intrusive Implementation of PR-BT
An immediate consequence is that positive-realness can be enforced non-intrusively on the ROM produced by I-PORK, using only samples of $G(s)$ at $\sigma_{i}$ and $G(\infty)$. This is implemented in Algorithm 2, named “Data-driven Positive-real Interpolation ($V$-type) (DD-PRI-V)”. The resulting ROM is simply a different state-space realization of the ROM defined in (44).
Algorithm 2 DD-PRI-V
Input: Interpolation points: $(\sigma_{1},\cdots,\sigma_{v})$; Data: $\big{(}H(\sigma_{1}),\cdots,H(\sigma_{v})\big{)}$ and $G(\infty)$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
1:  Compute $\hat{C}=C\hat{V}$ from the samples $H(\sigma_{i})$ via (5).
2:  Set $D=G(\infty)$ and $R_{\text{PR}}=D+D^{T}$.
3:  Set $S_{v}$ and $L_{v}$ as in (6).
4:  Compute $Q_{v}$ by solving the Lyapunov equation (9).
5:  Compute $T_{v}$ by solving the following Sylvester equation:$(S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}-Q_{v}^{-1}L_{v}^{T}R_{\text{PR}}^{-1}\hat{C})T_{v}-T_{v}S_{v}+Q_{v}^{-1}L_{v}^{T}R_{\text{PR}}^{-\frac{1}{2}}L_{v}=0$.
6:  Set $\hat{C}_{\text{PR}}=R_{\text{PR}}^{-\frac{1}{2}}\hat{C}T_{v}$.
7:  Update $Q_{v}$ by solving the Lyapunov equations (42).
8:  Set $\hat{B}=T_{v}Q_{v}^{-1}L_{v}^{T}R_{\text{PR}}^{\frac{1}{2}}$ and $\hat{A}=S_{v}-\hat{B}L_{v}$.
@@
An immediate consequence is that positive-realness can be enforced non-intrusively on the ROM produced by O-PORK, using only samples of $G(s)$ at $\mu_{i}$ and $G(\infty)$. This is implemented in Algorithm 3, named “Data-driven Positive-real Interpolation ($W$-type) (DD-PRI-W)”. The resulting ROM is simply a different state-space realization of the ROM defined in (55).
Algorithm 3 DD-PRI-W
Input: Interpolation points: $(\mu_{1},\cdots,\mu_{w})$; Data: $\big{(}H(\mu_{1}),\cdots,H(\mu_{w})\big{)}$ and $G(\infty)$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
1:  Compute $\hat{B}=\hat{W}^{*}B$ from the samples $H(\mu_{i})$ via (5).
2:  Set $D=G(\infty)$ and $R_{\text{PR}}=D+D^{T}$.
3:  Set $S_{w}$ and $L_{w}$ as in (6).
4:  Compute $P_{w}$ by solving the Lyapunov equations (10).
5:  Compute $T_{w}$ by solving the following Sylvester equation:$(S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}-\hat{B}R_{\text{PR}}^{-1}L_{w}^{T}P_{w}^{-1})^{*}T_{w}-T_{w}S_{w}^{*}+P_{w}^{-1}L_{w}R_{\text{PR}}^{-\frac{1}{2}}L_{w}^{T}=0$.
6:  Set $\hat{B}_{\text{PR}}=T_{w}^{*}\hat{B}R_{\text{PR}}^{-\frac{1}{2}}$.
7:  Update $P_{w}$ by solving the Lyapunov equations (53).
8:  Set $\hat{C}=R_{\text{PR}}^{\frac{1}{2}}L_{w}^{T}P_{w}^{-1}T_{w}^{*}$ and $\hat{A}=S_{w}-L_{w}\hat{C}$.
Similar to the ADI-based non-intrusive implementation of BT, the RADI-based non-intrusive implementation of PR-BT can be directly obtained using the results of Theorems 2.3 and 2.4. The pseudo-code for this implementation, named “Data-driven PR-BT (DD-PR-BT)”, is provided in Algorithm 4.
@@
Algorithm 4 DD-PR-BT
1:  Compute the Loewner quadruplet $(\hat{W}^{*}\hat{V},\hat{W}^{*}A\hat{V},\hat{W}^{*}B,C\hat{V})$ from the samples $H(\sigma_{i})$ and $H(\mu_{i})$ via (5).
@@
5:  Compute $T_{v}$ and $T_{w}$ by solving the following Sylvester equations:$(S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}-Q_{v}^{-1}L_{v}^{T}R_{\text{PR}}^{-1}C\hat{V})T_{v}-T_{v}S_{v}+Q_{v}^{-1}L_{v}^{T}R_{\text{PR}}^{-\frac{1}{2}}L_{v}=0$,$(S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}-\hat{W}^{*}BR_{\text{PR}}^{-1}L_{w}^{T}P_{w}^{-1})^{*}T_{w}-T_{w}S_{w}^{*}+P_{w}^{-1}L_{w}R_{\text{PR}}^{-\frac{1}{2}}L_{w}^{T}=0$.
@@
9:  Compute the following SVD:$\hat{L}_{q}^{*}T_{w}^{*}(\hat{W}^{*}\hat{V})T_{v}\hat{L}_{p}=\begin{bmatrix}\tilde{U}_{1}&\tilde{U}_{2}\end{bmatrix}\begin{bmatrix}\tilde{\Sigma}_{r}&0\\ 0&\tilde{\Sigma}_{n-r}\end{bmatrix}\begin{bmatrix}\tilde{V}_{1}^{*}\\ \tilde{V}_{2}^{*}\end{bmatrix}$.
@@ 2.3 RADI-based Non-intrusive Implementation of BR-BT
An immediate consequence of this observation is that bounded-realness can be enforced non-intrusively on the ROM produced by I-PORK using only samples of $G(s)$ at $\sigma_{i}$ and the sample $G(\infty)$. This is implemented in Algorithm 5, named “Data-driven Bounded-real Interpolation ($V$-type) (DD-BRI-V)”. The resulting ROM is simply a different state-space realization of the ROM defined in (64).
Algorithm 5 DD-BRI-V
Input: Interpolation points: $(\sigma_{1},\cdots,\sigma_{v})$; Data: $\big{(}H(\sigma_{1}),\cdots,H(\sigma_{v})\big{)}$ and $G(\infty)$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
1:  Compute $\hat{C}=C\hat{V}$ from the samples $H(\sigma_{i})$ via (5).
2:  Set $D=G(\infty)$ $R_{p}=I-DD^{T}$, and $R_{b}=I+D^{T}R_{p}^{-1}D$.
3:  Set $S_{v}$ and $L_{v}$ as in (6).
4:  Compute $Q_{v}$ by solving the Lyapunov equation (9).
5:  Compute $T_{v}$ by solving the following Sylvester equation:$(S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}+Q_{v}^{-1}L_{v}^{T}D^{T}R_{p}^{-1}C\hat{V})T_{v}-T_{v}S_{v}+Q_{v}^{-1}L_{v}^{T}R_{b}^{\frac{1}{2}}L_{v}=0$.
6:  Set $\hat{C}_{\text{BR}}^{\text{P}}=R_{p}^{-\frac{1}{2}}\hat{C}T_{v}$.
7:  Update $Q_{v}$ by solving the Lyapunov equations (62).
8:  Set $\hat{B}=T_{v}Q_{v}^{-1}L_{v}^{T}R_{b}^{-\frac{1}{2}}$ and $\hat{A}=S_{v}-\hat{B}L_{v}$.
@@
An immediate consequence is that bounded-realness can be enforced non-intrusively on the ROM produced by O-PORK, using only samples of $G(s)$ at $\mu_{i}$ and $G(\infty)$. This approach is implemented in Algorithm 6, named “Data-driven Bounded-real Interpolation ($W$-type) (DD-BRI-W)”. The resulting ROM represents a different state-space realization of the ROM defined in (70).
Algorithm 6 DD-BRI-W
Input: Interpolation points: $(\mu_{1},\cdots,\mu_{w})$; Data: $\big{(}H(\mu_{1}),\cdots,H(\mu_{w})\big{)}$ and $G(\infty)$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
1:  Compute $\hat{B}=\hat{W}^{*}B$ from the samples $H(\mu_{i})$ via (5).
2:  Set $D=G(\infty)$, $R_{q}=I-D^{T}D$, and $R_{c}=I+DR_{q}^{-1}D^{T}$.
3:  Set $S_{w}$ and $L_{w}$ as in (6).
4:  Compute $P_{w}$ by solving the Lyapunov equations (10).
5:  Compute $T_{w}$ by solving the following Sylvester equation:$(S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}+\hat{W}^{*}BR_{q}^{-1}D^{T}L_{w}^{T}P_{w}^{-1})^{*}T_{w}-T_{w}S_{w}^{*}+P_{w}^{-1}L_{w}R_{c}^{\frac{1}{2}}L_{w}^{T}=0$.
6:  Set $\hat{B}_{\text{PR}}=T_{w}^{*}\hat{B}R_{q}^{-\frac{1}{2}}$.
7:  Update $P_{w}$ by solving the Lyapunov equations (68).
8:  Set $\hat{C}=R_{c}^{-\frac{1}{2}}L_{w}^{T}P_{w}^{-1}T_{w}^{*}$ and $\hat{A}=S_{w}-L_{w}\hat{C}$.
Similar to the ADI-based non-intrusive implementation of BT, the RADI-based non-intrusive implementation of BR-BT can be directly obtained using the results of Theorems 2.5 and 2.6. The pseudo-code for this implementation, named “Data-driven BR-BT (DD-BR-BT)”, is provided in Algorithm 7.
@@
Algorithm 7 DD-BR-BT
1:  Compute the Loewner quadruplet $(\hat{W}^{*}\hat{V},\hat{W}^{*}A\hat{V},\hat{W}^{*}B,C\hat{V})$ from the samples $H(\sigma_{i})$ and $H(\mu_{i})$ via (5).
@@
5:  Compute $T_{v}$ and $T_{w}$ by solving the following Sylvester equations:$(S_{v}-Q_{v}^{-1}L_{v}^{T}L_{v}+Q_{v}^{-1}L_{v}^{T}D^{T}R_{p}^{-1}C\hat{V})T_{v}-T_{v}S_{v}+Q_{v}^{-1}L_{v}^{T}R_{b}^{\frac{1}{2}}L_{v}=0$,$(S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}+\hat{W}^{*}BR_{q}^{-1}D^{T}L_{w}^{T}P_{w}^{-1})^{*}T_{w}-T_{w}S_{w}^{*}+P_{w}^{-1}L_{w}R_{c}^{\frac{1}{2}}L_{w}^{T}=0$.
@@
9:  Compute the following SVD:$\hat{L}_{q}^{*}T_{w}^{*}(\hat{W}^{*}\hat{V})T_{v}\hat{L}_{p}=\begin{bmatrix}\tilde{U}_{1}&\tilde{U}_{2}\end{bmatrix}\begin{bmatrix}\tilde{\Sigma}_{r}&0\\ 0&\tilde{\Sigma}_{n-r}\end{bmatrix}\begin{bmatrix}\tilde{V}_{1}^{*}\\ \tilde{V}_{2}^{*}\end{bmatrix}$.
@@ 2.4 ADI-based Non-intrusive Implementation of SW-BT
An immediate consequence of this observation is that the minimum-phase property can be enforced on the ROM nonintrusively produced by O-PORK using only samples of $G(s)$ at $\mu_{i}$ and the sample $G(\infty)$. This is implemented in Algorithm 8, named “Data-driven Minimum-phase Interpolation (DD-MPI)”. The resulting ROM is simply a different state-space realization of the ROM defined in (75).
Algorithm 8 DD-MPI
Input: Interpolation points: $(\mu_{1},\cdots,\mu_{w})$; Data: $\big{(}H(\mu_{1}),\cdots,H(\mu_{w})\big{)}$ and $G(\infty)$.
Output: ROM: $(\hat{A},\hat{B},\hat{C},D)$
1:  Compute $\hat{B}=\hat{W}^{*}B$ from the samples $H(\mu_{i})$ via (5) and set $D=G(\infty)$..
2:  Set $S_{w}$ and $L_{w}$ as in (6).
3:  Compute $P_{w}$ by solving the Lyapunov equations (10).
4:  Compute $T_{w}$ by solving the following Sylvester equation:$\big{(}S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}-\hat{B}D^{-1}L_{w}^{T}P_{w}^{-1}\big{)}^{*}T_{w}-T_{w}S_{w}^{*}+P_{w}^{-1}L_{w}D^{-T}L_{w}^{T}=0$.
5:  Set $\hat{C}=DL_{w}^{T}P_{w}^{-1}T_{w}^{*}$ and $\hat{A}=S_{w}-L_{w}\hat{C}$.
Similar to the ADI-based nonintrusive implementation of BT, the ADI-based nonintrusive implementation of SW-BT can be derived directly using the results of Theorem 2.7. The pseudo-code for this implementation, named “Data-driven SW-BT (DD-SW-BT)”, is provided in Algorithm 9.
@@
Algorithm 9 DD-SW-BT
1:  Compute the Loewner quadruplet $(\hat{W}^{*}\hat{V},\hat{W}^{*}A\hat{V},\hat{W}^{*}B,C\hat{V})$ from the samples $H(\sigma_{i})$ and $H(\mu_{i})$ via (5) and set $D=G(\infty)$.
@@
6:  Compute the following SVD:$\hat{L}_{q}^{*}T_{w}^{*}(\hat{W}^{*}\hat{V})\hat{L}_{p}=\begin{bmatrix}\tilde{U}_{1}&\tilde{U}_{2}\end{bmatrix}\begin{bmatrix}\tilde{\Sigma}_{r}&0\\ 0&\tilde{\Sigma}_{n-r}\end{bmatrix}\begin{bmatrix}\tilde{V}_{1}^{*}\\ \tilde{V}_{2}^{*}\end{bmatrix}$.
@@ 2.5 ADI/RADI-based Non-intrusive Implementation of BST
where $\hat{B}_{\text{S}}=W_{\text{S}}^{*}\tilde{B}_{\text{S}}$.
@@
$\big{(}S_{w}-L_{w}\zeta-[(\hat{W}^{*}\hat{V})Q_{v}^{-1}(C\hat{V})^{*}+(\hat{W}^{*}B)D^{T}]R_{\text{S}}^{-1}\zeta\big{)}^{*}T_{w}-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{S}}^{-\frac{1}{2}}L_{w}^{T}=0.$
Similar to the ADI-based nonintrusive implementation of BT, the ADI- and RADI-based nonintrusive implementation of BST can be obtained readily. The pseudo-code for this implementation, named “Data-driven BST (DD-BST)”, is provided in Algorithm 10.
@@
Algorithm 10 DD-BST
1:  Compute the Loewner quadruplet $(\hat{W}^{*}\hat{V},\hat{W}^{*}A\hat{V},\hat{W}^{*}B,C\hat{V})$ from the samples $H(\sigma_{i})$ and $H(\mu_{i})$ via (5).
@@
5:  Compute $T_{w}$ by solving the following Sylvester equation:$(S_{w}-L_{w}L_{w}^{T}P_{w}^{-1}-[(\hat{W}^{*}\hat{V})Q_{v}^{-1}(C\hat{V})^{*}+(\hat{W}^{*}B)D^{T}]R_{\text{S}}^{-1}L_{w}^{T}P_{w}^{-1})^{*}T_{w}-T_{w}S_{w}^{*}+P_{w}^{-1}L_{w}R_{\text{S}}^{-\frac{1}{2}}L_{w}^{T}=0$.
@@
9:  Compute the following SVD:$\hat{L}_{q}^{*}T_{w}^{*}(\hat{W}^{*}\hat{V})\hat{L}_{p}=\begin{bmatrix}\tilde{U}_{1}&\tilde{U}_{2}\end{bmatrix}\begin{bmatrix}\tilde{\Sigma}_{r}&0\\ 0&\tilde{\Sigma}_{n-r}\end{bmatrix}\begin{bmatrix}\tilde{V}_{1}^{*}\\ \tilde{V}_{2}^{*}\end{bmatrix}$.
@@ 3 Practical Considerations in Implementation
The theoretical foundations of the ADI [19] and RADI [4] methods are well established. Their performance in approximating Lyapunov and Riccati equations is also well documented. However, the implementation presented in this paper involves inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$, which can cause numerical problems as the number of samples used for implementation increases. This is because the solution of Lyapunov equations is generally low-rank, enabling low-rank approximation via the low-rank ADI method [19] in the first place. Thus, as the number of samples used for implementation increases, the matrices $Q_{v}$ and $P_{w}$ start losing rank. Note that these matrices do not appear in the original algorithms. In data-driven modeling, one typically uses as many samples as available to ensure maximum accuracy. As discussed in [2], the inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$ can be computed conveniently when the interpolation points are lightly damped, i.e.,
@@
wherein $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$. If we select the damping coefficients $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, this means we are sampling $G(s)$ in close proximity to the $j\omega$-axis, which is also practically reasonable for capturing the frequency response of $G(s)$. When the interpolation points are lightly damped (i.e., $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$), $Q_{v}$, $P_{w}$, $Q_{v}^{-1}$, and $P_{w}^{-1}$ become diagonally dominant. If $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, the off-diagonal elements are nearly zero. For details, refer to [22]. Essentially, this means that each diagonal element of $Q_{v}^{-1}$ and $P_{w}^{-1}$ depends solely on its respective pair $(\sigma_{i},G(\sigma_{i}))$ or $(\mu_{i},G(\mu_{i}))$, while the dependence on other samples or sampling points becomes negligible.
Assuming that the interpolation points are lightly damped, we can derive the expressions for the diagonal entries of $Q_{v}^{-1}$ and $P_{w}^{-1}$ as follows:
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\sigma_{1})[I_{m}+H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2Real(\sigma_{v})[I_{m}+H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big{)},$ (81)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\mu_{1})[I_{p}+H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2Real(\mu_{w})[I_{p}+H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big{)}.$ (82)
For $\mathcal{H}_{\infty}$-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\sigma_{1})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2Real(\sigma_{v})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big{)},$ (83)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\mu_{1})[I_{p}+(1-\gamma^{-2})H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2Real(\mu_{w})[I_{p}+(1-\gamma^{-2})H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big{)}.$ (84)
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\sigma_{1})[I_{m}-t_{v,1}^{*}H^{*}(\sigma_{1})R_{\text{PR}}^{-1}H(\sigma_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2Real(\sigma_{v})[I_{m}-t_{v,v}^{*}H^{*}(\sigma_{v})R_{\text{PR}}^{-1}H(\sigma_{v})t_{v,v}]^{-1}\big{)},$ (85)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\mu_{1})[I_{p}-t_{w,1}^{*}H(\mu_{1})R_{\text{PR}}^{-1}H^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2Real(\mu_{w})[I_{p}-t_{w,w}^{*}H(\mu_{w})R_{\text{PR}}^{-1}H^{*}(\mu_{w})t_{w,w}]^{-1}\big{)},$ (86)
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\sigma_{1})[I_{m}-t_{v,1}^{*}H^{*}(\sigma_{1})R_{p}^{-1}H(\sigma_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2Real(\sigma_{v})[I_{m}+t_{v,v}^{*}H^{*}(\sigma_{v})R_{p}^{-1}H(\sigma_{v})t_{v,v}]^{-1}\big{)},$ (91)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\mu_{1})[I_{p}-t_{w,1}^{*}H(\mu_{1})R_{q}^{-1}H^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2Real(\mu_{w})[I_{p}-t_{w,w}^{*}H(\mu_{w})R_{q}^{-1}H^{*}(\mu_{w})t_{w,w}]^{-1}\big{)},$ (92)
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\sigma_{1})I_{m},\cdots,2Real(\sigma_{v})I_{m}\big{)},$ (97)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big{(}2Real(\mu_{1})I_{p},\cdots,2Real(\mu_{w})I_{p}\big{)},$ (98)

### v2

@@ 3.2 RADI-based Non-intrusive Implementation of PR-BT
An immediate consequence is that positive-realness can be enforced non-intrusively on the ROM produced by I-PORK, using only samples of $G(s)$ at $\sigma_{i}$ and $G(\infty)$.
@@
An immediate consequence is that positive-realness can be enforced non-intrusively on the ROM produced by O-PORK, using only samples of $G(s)$ at $\mu_{i}$ and $G(\infty)$.
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of PR-BT can be directly obtained using the results of Theorems 3.3 and 3.4. The pseudo-code for this implementation, named “Data-driven PR-BT (DD-PR-BT)”, is provided in Algorithm 2.
@@
Algorithm 2 $D$-PR-BT
[…]
@@
[…]
@@
[…]
@@ 3.3 RADI-based Non-intrusive Implementation of BR-BT
The RADI-based approximation of $P_{\text{BR}}$ is given by $V_{\text{BR}}Q_{v}^{-1}V_{\text{BR}}^{*}$.
@@
An immediate consequence of this observation is that bounded-realness can be enforced non-intrusively on the ROM produced by I-PORK using only samples of $G(s)$ at $\sigma_{i}$ and the sample $G(\infty)$.
@@
An immediate consequence is that bounded-realness can be enforced non-intrusively on the ROM produced by O-PORK, using only samples of $G(s)$ at $\mu_{i}$ and $G(\infty)$.
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of BR-BT can be directly obtained using the results of Theorems 3.5 and 3.6. The pseudo-code for this implementation, named “Data-driven BR-BT (DD-BR-BT)”, is provided in Algorithm 3.
@@
Algorithm 3 $D$-BR-BT
[…]
@@
[…]
@@
[…]
@@ 3.4 ADI-based Non-intrusive Implementation of SW-BT
An immediate consequence of this observation is that the minimum-phase property can be enforced on the ROM nonintrusively produced by O-PORK using only samples of $G(s)$ at $\mu_{i}$ and the sample $G(\infty)$.
Similar to DD-ADI-BT [2], the ADI-based nonintrusive implementation of SW-BT can be derived directly using the results of Theorem 3.7. The pseudo-code for this implementation, named “Data-driven SW-BT (DD-SW-BT)”, is provided in Algorithm 4.
@@
Algorithm 4 $D$-SW-BT
[…]
@@
[…]
@@ 3.5 ADI/RADI-based Non-intrusive Implementation of BST
where $\hat{B}_{\text{S}}=W_{\text{S}}^{*}\tilde{B}_{\text{S}}$. The RADI-based approximation of $\tilde{Q}_{\text{S}}$ is given by $W_{\text{S}}P_{w}^{-1}W_{\text{S}}^{*}$.
@@
$\big(S_{w}-L_{w}\zeta-[(\hat{W}^{*}\hat{V})Q_{v}^{-1}(C\hat{V})^{*}+(\hat{W}^{*}B)D^{T}]$ $R_{\text{S}}^{-1}\zeta\big)^{*}T_{w}$
$-T_{w}S_{w}^{*}+\zeta^{*}R_{\text{S}}^{-\frac{1}{2}}L_{w}^{T}=0.$
Similar to DD-ADI-BT [2], the ADI- and RADI-based nonintrusive implementation of BST can be obtained readily. The pseudo-code for this implementation, named “Data-driven BST (DD-BST)”, is provided in Algorithm 5.
@@
Algorithm 5 $D$-BST
[…]
@@
[…]
@@
[…]
@@ 4 Practical Considerations in Implementation
The theoretical foundations of the ADI [16] and RADI [4] methods are well established. Their performance in approximating Lyapunov and Riccati equations is also well documented. However, the implementation presented in this paper involves inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$, which can cause numerical problems as the number of samples used for implementation increases. This is because the solution of Lyapunov equations is generally low-rank, enabling low-rank approximation via the low-rank ADI method [16] in the first place. Thus, as the number of samples used for implementation increases, the matrices $Q_{v}$ and $P_{w}$ start losing rank. Note that these matrices do not appear in the original algorithms. In data-driven modeling, one typically uses as many samples as available to ensure maximum accuracy. As discussed in [2], the inverses $Q_{v}^{-1}$ and $P_{w}^{-1}$ in DD-ADI-BT can be computed conveniently when the interpolation points are lightly damped, i.e.,
@@
wherein $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$. If we select the damping coefficients $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, this means we are sampling $G(s)$ in close proximity to the $j\omega$-axis, which is also practically reasonable for capturing the frequency response of $G(s)$. When the interpolation points are lightly damped (i.e., $\zeta_{i,\sigma}\ll 1$ and $\zeta_{i,\mu}\ll 1$), the respective matrices $Q_{v}$, $P_{w}$, $Q_{v}^{-1}$, $P_{w}^{-1}$, $T_{v}$, and $T_{w}$ become diagonally dominant due to the modal form of the matrices $S_{v}$ and $S_{w}$ that appear in their associated Lyapunov, Sylvester, and Riccati equations. If $\zeta_{i,\sigma}\approx 0$ and $\zeta_{i,\mu}\approx 0$, the off-diagonal elements are nearly zero. For details and analytical formulas for computing the diagonal elements of the Lyapunov, Sylvester, and Riccati equations in this case, see [40]. Essentially, this means that each diagonal element of $Q_{v}^{-1}$ and $P_{w}^{-1}$ depends solely on its respective pair $(\sigma_{i},G(\sigma_{i}))$ or $(\mu_{i},G(\mu_{i}))$, while the dependence on other samples or sampling points becomes negligible.
Assuming that the interpolation points are lightly damped, the expressions for the diagonal entries of $Q_{v}^{-1}$ and $P_{w}^{-1}$ can be derived using the analytical formulas in [40] as follows:
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}+H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\text{Real}(\sigma_{v})[I_{m}+H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big),$ (79)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}+H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 113.81102pt2\text{Real}(\mu_{w})[I_{p}+H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big).$ (80)
For $D$-$\mathcal{H}_{\infty}$-BT:
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{1})H(\sigma_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\text{Real}(\sigma_{v})[I_{m}+(1-\gamma^{-2})H^{*}(\sigma_{v})H(\sigma_{v})]^{-1}\big),$ (81)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}+(1-\gamma^{-2})H(\mu_{1})H^{*}(\mu_{1})]^{-1},\cdots,$
$\hskip 85.35826pt2\text{Real}(\mu_{w})[I_{p}+(1-\gamma^{-2})H(\mu_{w})H^{*}(\mu_{w})]^{-1}\big).$ (82)
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}-t_{v,1}^{*}H^{*}(\sigma_{1})R_{\text{PR}}^{-1}H(\sigma_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\sigma_{v})[I_{m}-t_{v,v}^{*}H^{*}(\sigma_{v})R_{\text{PR}}^{-1}H(\sigma_{v})t_{v,v}]^{-1}\big),$ (83)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}-t_{w,1}^{*}H(\mu_{1})R_{\text{PR}}^{-1}H^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\mu_{w})[I_{p}-t_{w,w}^{*}H(\mu_{w})R_{\text{PR}}^{-1}H^{*}(\mu_{w})t_{w,w}]^{-1}\big),$ (84)
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})[I_{m}-t_{v,1}^{*}H^{*}(\sigma_{1})R_{p}^{-1}H(\sigma_{1})t_{v,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\sigma_{v})[I_{m}+t_{v,v}^{*}H^{*}(\sigma_{v})R_{p}^{-1}H(\sigma_{v})t_{v,v}]^{-1}\big),$ (89)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})[I_{p}-t_{w,1}^{*}H(\mu_{1})R_{q}^{-1}H^{*}(\mu_{1})t_{w,1}]^{-1},\cdots,$
$\hskip 56.9055pt2\text{Real}(\mu_{w})[I_{p}-t_{w,w}^{*}H(\mu_{w})R_{q}^{-1}H^{*}(\mu_{w})t_{w,w}]^{-1}\big),$ (90)
@@
$Q_{v}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\sigma_{1})I_{m},\cdots,2\text{Real}(\sigma_{v})I_{m}\big),$ (95)
$P_{w}^{-1}$ $\approx\text{blkdiag}\big(2\text{Real}(\mu_{1})I_{p},\cdots,2\text{Real}(\mu_{w})I_{p}\big),$ (96)